Shylls, Inc. determines that its marginal revenue per day is given by R'|t|=100et and that R(0)=0, where R(t) is the total accumulated revenue, in dollars, on day t. The company's marginal cost, per day, is given by C'|t|=100-0.2t with C(0)=0, where C(t) is the total accumulated cost, in dollars, on day t. What is the average daily profit for the first 10 days (t=0 to t=10)?
Answer #1

1 Sol? R(t) = 100et R(t) = 5100edt +C = 100e e. RO) = 0 + 0 = 100000)+c = c =-100. R(t) = 100e -100. C(B) = 100 -0.27 = C(